# BIFROST: A First-Principles Model of Polarization Mode Dispersion in Optical Fiber

### v1

@@ I Introduction
In many of these applications, the polarization of the guided light is a critical consideration. Due to birefringences in the fiber, the polarization varies in time, a phenomenon known as polarization mode dispersion (PMD) [12, 13]. PMD is an enabling phenomenon for some applications and is detrimental for others. In interferometric fiber sensors [4, 14, 15], PMD causes polarization fading [16], decreasing the sensitivity, while in fiber Bragg grating sensors [17], PMD variation causes drifts in the resonance condition [18]. Conversely, birefringence enables sensing of directional forces [5, 19, 6, 7]. Additionally, interferometry techniques to improve the resolution and sensitivity of optical telescopes involve fiber and its polarization properties [20, 21]. Finally, in the context of quantum networks [22], addressing the limits due to birefringences in fiber [10, 23, 24] is an active area of research [25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40].
@@
In this work, we construct a first-principles model of PMD in optical fibers. The model is implemented in an open-source Python library [45] known as the Birefringence in Fiber: Research and Optical Simulation Toolkit (BIFROST). This toolkit is a bridge between prior literature and the needs of current application domains. (In this work, we will use “BIFROST” to refer to both the Python implementation of the model and to the model itself.) We begin by reviewing the history of optical fibers relevant for our work. Then we describe the top-level model of BIFROST, review the physics included in the model, and enumerate the data sources used in the numerical calculation of fiber parameters. Then we compare prior work to simulations from BIFROST to validate our model. Finally, we perform a brief example investigation with BIFROST to demonstrate its utility, and discuss future directions.
@@ II A Brief History of Optical Fibers
Among the earliest live communication signals sent via optical fiber was telephone traffic at hundreds of Mb/skm, in 1977 [46]. Already by this time, birefringence in fiber was identified as a possible limit on data rates [47, 48, 46]. The reason for this is shown in Fig. 1: in on-off intensity schemes with no polarization control, birefringence causes the parts of the light along the fast and slow axes to propagate at different group velocities, resulting in pulse broadening [49]. If a pulse broadens too much, it can interfere with adjacent data frames, limiting the data rate [46]. Initial measurements in short (- m) fibers [47], when extrapolated linearly with fiber length, suggested a PMD-limited data rate of Gb/skm. Subsequent direct measurements on long ( km) fiber links [50] and a flurry of theoretical activity [e.g. 48, 51, 52] showed that PMD actually scales as the square root of the length. This increased the estimated limit due to PMD to tens of Tb/skm [50].
@@
An improvement on the situation, first introduced in 2006 and widely adopted since, is the so-called hinge model [61]. This model, shown in Fig. 2, describes buried fibers as a set of long segments alternating with short lengths called “hinges.” The long segments correspond to the buried parts of the fiber, and are modeled as stable in time because underground environmental variations are small. Hinges, on the other hand, correspond to above-ground segments, e.g. at repeater stations, data closets, and points of presence. These hinges, being above ground, experience environmental fluctuations and resulting birefringence changes; in the hinge model, all of the time variation is in the hinges, which has been traditionally modeled either with a predetermined rotation function [61] or with uniform random walks on the Poincaré sphere [62, 44] with step size chosen to fit to observed statistics. In these implementations, the long time-invariant segments are specified by a fixed total length and a desired PMD, and are subdivided by arbitrary fixed rotations to reproduce observed statistical distributions of PMD. The hinge model has had some success in reproducing the observed output statistics of PMD in buried fibers [61, 44], but it is limited: there is no connection to physical parameters such as operating temperature or external stress, making it impossible to explore how a fiber link responds to environmental changes.
@@ III The Top-Level Model of BIFROST
In BIFROST, we adapt the hinge model to include physically motivated (and physically constrained) parameters (Fig. 2). We model fiber as a step-index cylindrical waveguide, deriving refractive indices from material properties and environmental variables such as temperature. Birefringences from multiple physical mechanisms are included. Long, stable segments are specified by fiber geometry, core/cladding glass types, bend radii and twist rates where appropriate, and operating temperature. Each hinge is modeled as a set of fiber paddles, chosen because they are common experimental tools for manipulating polarization. The paddles are defined by a radius of curvature, number of fiber turns, and paddle angle. Additionally, we model spinning by inserting arbitrary Jones matrix rotations in the long fiber segments, with spacing determined by typical fiber spinning periods.
Thus, rather than specifying a desired hinge or segment PMD, BIFROST uses physical parameters to calculate an expected PMD. Grounding the calculations in physical birefringence mechanisms results in a model that can describe the changes in the output polarization due to temperature, wavelength variation, and more, making BIFROST a powerful predictive tool.
@@ IV.1 Modes of a Cylindrical Waveguide
is the normalized frequency parameter [unitless]. The fiber is single-mode when (of course, this means only one spatial mode propagates in the fiber; there are two orthogonally polarized modes).
@@ IV.2 Chromatic Dispersion
where [unitless] are resonance strengths and [m] are the resonance wavelengths, and the sum runs over as many resonances as are practically measurable or necessary for the intended application. In fused silica, usually either three [67, 68, 66] or two [69, 70] resonances are used. In both cases, at least one resonance in the deep UV captures the effect of photons promoting electrons to higher energy bands of the material, while at least one resonance in the infrared captures excitation of vibrational modes of the glass [69]. The refractive index of fused silica is shown in the top panel of Fig. 3, indicating both a UV resonance at the short wavelengths and the decrease in due to lattice absorption at long wavelengths.
@@ IV.3 Sources of Birefringence
When the core of the fiber is not circular, the boundary conditions that determine the propagation constants are different. Letting the ellipse be oriented with its semimajor and semiminor axes along the - and -axes (explicitly, ), we obtain two effective parameters . We also define the eccentricity [unitless] 111If , all of the expressions in this section flip .. For small core eccentricities, we can use the Gaussian approximation to derive the resulting birefringence [63]. For and to second order in , it is [63, 74, 75]
@@
Here is the operating temperature of the fiber, [unitless], and [unitless] are the photoelastic constants of the core, and are the thermal expansion coefficients for the core and cladding, and [unitless] is Poisson’s ratio for the core.
@@
where [unitless] is a commonly defined combination of material constants loosely called the “strain-optic coefficient.”
@@
When a fiber is twisted, a shear stress is introduced, which couples the transverse field of one mode to the longitudinal field of another. (Because we work in the weakly guiding limit, we have largely been ignoring the longitudinal field components, but they are critical here. In a typical Corning SMF-28e fiber at 1550 nm, the longitudinal field amplitude is a few percent of the transverse field amplitude.) The transverse and longitudinal fields are out of phase, so the coupling results in a circular birefringence. The birefringence is proportional to the twist rate [rad/m], and can be written [12]
@@ IV.5 Modeling Spun Fiber
where the Pauli matrices and is the identity matrix. The family of Jones matrices sampled in this way implement uniform rotations on the Poincaré sphere [62, 44].
@@ IV.6 Differential Group Delay
such that the total delay between these two modes is exactly from Eqn. 21. The eigenmodes in this case are and (where T is the transpose), which are sensible as these states are aligned with the fast and slow axes of the fiber. The eigenmodes are called principal states of polarization (PSPs); they do not vary to first order with frequency, and the time delay between them is exactly the DGD of the time-domain picture.
@@
The DGD is a single scalar value, and therefore insufficient to describe the full PTF. Nevertheless, the DGD is still helpful as a heuristic metric of polarization rotation in the fiber. To see this, let us consider putting a temporally long pulse through each of three different fibers and consider the consequence of changing the fiber temperature. (1) If the fiber has no birefringence, then there is no PMD, and therefore zero DGD, and also the SOP remains the same at the output as at the input. The temperature change has no effect on the output state. (2) Conversely, in a long segment of fiber with constant birefringence, the delay between the two polarization components is allowed to build up over a long distance, so the total DGD is large. At the same time the SOP is also rotating along the fiber length, so the output SOP is very sensitive to changes in temperature. (3) Finally, for a long spun fiber, the spinning scrambles the birefringence axes; thus no two polarization modes are allowed to build up large relative delays, and the total DGD is small (in comparison to an unspun fiber of the same length). As the light traverses the fiber, the SOP does a random walk on the Poincaré sphere, rotating around one axis for a short length, then another axis for a short length, and so on. Such a random walk is hard to track, but the polarization does not wander as quickly as a function of length down the fiber. When the temperature is changed, each of these random walk steps is lengthened or shortened slightly, making it challenging to predict how the output SOP changes, but since the changes of length are in arbitrary directions, the total change to the SOP is smaller than for the unspun fiber of the same length.
@@ V Calculation Data and Methods
Our aim in this work is to model typical commercial telecommunications fiber links. Much of this fiber is predominantly silica-germania glass [82], i.e. pure silica (SiO2) for the cladding, and silica doped with a small amount of germania (GeO2) for the core, so we model these materials in our work. The material properties of bulk silica and bulk germania have been well-studied, giving us a wealth of data to use in BIFROST. To specify the effects of germania doping, we use the mole percent (%mol) (the fraction of molecules in the glass that are germania [82]) to specify the concentration, and we use a linear additive model for each of the relevant fiber properties. That is, for a given fiber property which has a value for pure silica and for pure germania, the value we use for silica with molar fraction of germania doping is
@@
To calculate indices of refraction, we use three-term Sellmeier equations for both silica and germania. For silica, we use the Sellmeier coefficients from Ref. [66]. The authors measured the refractive index of Corning 7980 fused silica in the ranges and , and fit the data to a set of three Sellmeier resonances with quartic temperature dependence. Thus the Sellmeier equation for silica takes into account the temperature dependence of the refractive index. For germania, we use the data from Refs. [87, 88], measured at an average of 24∘C over the wavelength range from 365 nm to 4.3 m. The authors do not know of any literature on the temperature-dependence of the coefficients of a three-term Sellmeier model of germania (though see Ref. [89] for the two-term case), but Ref. [82] gives an expression, fitted from measurements, for the thermo-optic coefficient of bulk germania glass in the range 200-300 K at 1550 nm. We therefore calculate the refractive index of germania at a temperature by calculating the expected value at 24∘ from Refs. [87, 88] and adding the expected change in the refractive index using the thermo-optic coefficient expression of Ref. [82].
@@ VI Validation of BIFROST
Having assembled a model of PMD in optical fibers, we need to verify that this model reproduces experimentally observed results. In some sense there is a wealth of data against which we can compare BIFROST results, since many works have made measurements of PMD in optical fibers (e.g. [91, 92, 93, 94] from the telecommunications engineering literature and [23, 24, 10] from the more recent quantum network literature). But in a more rigorous sense, the necessary experimental data for validation is scarce. BIFROST brings together many different “submodels,” including the model of the propagation constant of the fundamental modes Eqn. 2, the models of different birefringence mechanisms Eqns. 6, 8, 10, and 11, and the linear additive model for germania doping Eqn. 25 described in the last section. Ideally, each of these submodels should be compared to experiments which separate out the effects of the submodel under test, which is challenging experimentally. Such experiments, to the extent it is even possible to perform them, are rare in the literature. For this reason, validation of BIFROST is ongoing, and we are explicit about what we have examined thus far. We divide validation methodologies into five categories.
@@
In addition to verifying intrinsic fiber properties, it is important to verify the models of birefringence used in BIFROST. While all the equations we used are conventional models, each of them has limits; for instance, Eqn. 6 only works in the limit of small ellipticity, and Eqn. 10 may fail when the bend radii are small (i.e. when the bend is sharp). Thus it would be useful to check that each model works within reasonable regimes and test the degree of their failure in extreme regimes. To test such models requires very simple fiber systems whose properties we know well.
As an example of this type of validation, we consider fiber paddle systems such as ThorLabs FPC563. ThorLabs specifies [97] the total retardance of a single paddle given the paddle radius and the number of turns of fiber on the paddle. For instance, the manual specifies that six turns of fiber with cladding diameter of m wound onto a 56 mm diameter paddle should yield a retardance of half a wave at nm. Fig. 5 confirms that BIFROST reproduces this result. We simulated a single paddle with the above parameters, and we specified input light with linear polarization at 1400 nm. The output polarization is shown on the Poincaré sphere as the paddle angle is varied, equivalent to turning a waveplate in free space 444We also specify in the simulation that there is 2 cm of straight fiber length at the input of the paddle, which twists as the paddle is rotated; we “measure” the polarization right out of the loop. The twist rotates the birefringence axes relative to the input, which makes it physically similar to rotating a waveplate in free space.. The resulting polarization rotation is almost exactly that expected of a turning half-wave plate.
@@
As was discussed previously, there have been a number of previous attempts to model PMD in long fibers; some of these works ([e.g. 43, 61, 74, 75]) report on the specifics of their modeling choices and discuss specific birefringence mechanisms. In general, these works are not recent, and in the rare instances these works include comparison to experiment, the experiment is not well-documented. Nevertheless, comparisons between BIFROST and these previous results may show that BIFROST can reproduce previous computational results.
@@
In both cases, qualitative agreement is achieved: the birefringences due to core ellipticity and asymmetric thermal stress show similar shapes as a function of the normalized frequency , and the histograms from BIFROST fit well to Maxwellian curves just like the ones from Ref. [43]. However, there is quantitative disagreement of evident in both figures. The discrepancy appears separately for both the core ellipticity and the asymmetric thermal stress values. For the core ellipticity, Ref. [43] uses a different calculation method to approximate the birefringence due to core ellipticity, so we are perhaps unsurprised that there is some quantitative difference. (The method of Ref. [43] may be more accurate than our approximation Eqn. 6, and may be incorporated into BIFROST in the future.) For the asymmetric thermal stress, the difference might be due to the values of refractive indices used ( in Ref. [43] versus 1.44508 from BIFROST) or the calculation method for . The factor in Eqn. 8 is extremely numerically sensitive, because is the small difference of the two large numbers and . Indeed, quantitatively matching BIFROST’s calculations to those of Ref. [43] can be accomplished by manually increasing by just 0.05%. It is possible that the analytical approximation in Eqn. 2 needs to be revisited to aim for a more accurate result (where “accurate” here is relative to a numerical evaluation of the modes of an elliptical core).
In short, the comparisons of Fig. 6 point to good qualitative agreement, while the quantitative discrepancies are attributable to known differences in computational methods. This adds to the validity of BIFROST.
@@
Lastly, we may try to use BIFROST to directly simulate fibers whose PMD values are measured in experiments. Candidates for this type of comparison have been identified [103, 91, 92, 23, 24]. In all of these cases, the fiber is not so well-reported that the simulation parameters to be used in BIFROST are obvious from the manuscripts. However, BIFROST’s ability to analyze statistical ensembles allows us to explore the available parameter space and understand the parameter regimes under which BIFROST achieves qualitative and quantitative agreement with observations.

### v2

@@ I Introduction
In many of these applications, the polarization of the guided light is a critical consideration. Due to birefringences in the fiber, the polarization varies in time, a phenomenon known as polarization mode dispersion (PMD) [90, 49]. PMD is an enabling phenomenon for some applications and is detrimental for others. In interferometric fiber sensors [63, 55, 65], PMD causes polarization fading [115], decreasing the sensitivity, while in fiber Bragg grating sensors [43], PMD variation causes drifts in the resonance condition [104]. Conversely, birefringence enables sensing of directional forces [26, 15, 69, 78]. Additionally, interferometry techniques to improve the resolution and sensitivity of optical telescopes involve fiber polarization properties [50, 80]. Finally, in the context of quantum networks [7], addressing the limits due to fiber birefringence is an active area of research [30, 77, 64, 35, 98, 67, 72, 71, 85, 119, 110, 1, 101, 28, 108, 97, 113, 38, 114].
@@
In this work, we construct a first-principles model of PMD in optical fibers, in which all parts of the fiber and its behavior in time are modeled by fiber geometries, materials, and stresses. The model is implemented in an open-source Python library [10] known as the Birefringence in Fiber: Research and Optical Simulation Toolkit (BIFROST). This toolkit is a bridge between prior literature and the needs of current application domains. (In this work, we will use “BIFROST” to refer to both the Python implementation of the model and to the model itself.) We begin by reviewing the history of optical fibers relevant for our work. Then we describe the top-level model of BIFROST, review the physics included in the model, and enumerate the data sources used in the numerical calculation of fiber parameters. Then we compare prior work to simulations from BIFROST to validate our model. Finally, we perform a brief example investigation with BIFROST to demonstrate its utility, and discuss future directions.
@@ II A Brief History of Optical Fibers
Among the earliest live communication signals sent via optical fiber was telephone traffic at hundreds of Mb/skm, in 1977 [112]. Already by this time, birefringence in fiber was identified as a possible limit on data rates [88, 86, 112]. The reason for this is shown in Fig. 1: in on-off intensity schemes with no polarization control, birefringence causes the parts of the light along the fast and slow axes to propagate at different group velocities, resulting in pulse broadening [117]. If a pulse broadens too much, it can interfere with adjacent data frames, limiting the data rate [112]. Initial measurements in short (- m) fibers [88], when extrapolated linearly with fiber length, suggested a PMD-limited data rate of Gb/skm. Subsequent direct measurements on long ( km) fiber links [79] and a flurry of theoretical activity [e.g. 86, 41, 46] showed that PMD actually scales as the square root of the length. This increased the estimated limit due to PMD to tens of Tb/skm [79].
@@
An improvement on the situation, first introduced in 2006 and widely adopted since, is the so-called hinge model [20]. This model, shown in Fig. 2, describes buried fibers as a set of long segments alternating with short lengths called “hinges.” The long segments correspond to the buried parts of the fiber, and are modeled as stable in time because underground environmental variations are small. Hinges, on the other hand, correspond to above-ground segments, e.g. at repeater stations, data closets, and points of presence, which experience environmental fluctuations and resulting birefringence changes. In the hinge model, all of the time variation is in the hinges; this variation in time has been traditionally modeled either with a predetermined rotation function [20] or with uniform random walks on the Poincaré sphere [32, 31] with step size chosen to fit to observed statistics. In these implementations, the long time-invariant segments are specified by a fixed total length and a desired PMD, and are subdivided by arbitrary fixed rotations to reproduce observed statistical distributions of PMD. The hinge model has had some success in reproducing the observed output statistics of PMD in buried fibers [20, 31], but it is limited: there is no connection to physical parameters such as operating temperature or external stress, making it impossible to explore how a fiber link responds to environmental changes.
@@ III The Top-Level Model of BIFROST
In BIFROST, we adapt the hinge model to include physically motivated (and physically constrained) parameters (Fig. 2). We model fiber as a step-index cylindrical waveguide, deriving refractive indices from material properties and environmental variables such as temperature. Birefringences from multiple physical mechanisms are included. Long, stable segments are specified by fiber geometry, core/cladding glass types, bend radii and twist rates where appropriate, and operating temperature. Each hinge is modeled as a set of fiber paddles, chosen because they are common experimental tools for manipulating polarization. The paddles are defined by a radius of curvature, number of fiber turns, and paddle angle. Variation in time must be modeled according to how environmental changes, like temperature fluctuations, affect the physical parameters of the paddles. Additionally, we model spinning by inserting arbitrary Jones matrix rotations in the long fiber segments, with spacing determined by typical fiber spinning periods.
Thus, rather than specifying a desired hinge or segment PMD, BIFROST uses physical parameters to calculate an expected PMD. Unlike previous models, both our segments and our hinges are fully specified by physical parameters. Grounding the calculations in physical birefringence mechanisms results in a model that can describe the changes in the output polarization due to temperature, wavelength variation, and more, making BIFROST a powerful predictive tool.
@@ IV.1 Modes of a Cylindrical Waveguide
is the normalized (dimensionless) frequency parameter. The fiber is single-mode when (of course, this means only one spatial mode propagates in the fiber; there are two orthogonally polarized modes).
@@ IV.2 Chromatic Dispersion
where [dimensionless] are resonance strengths and [m] are the resonance wavelengths, and the sum runs over as many resonances as are practically measurable or necessary for the intended application. In fused silica, usually either three [74, 76, 66] or two [44, 3] resonances are used. In both cases, at least one resonance in the deep UV captures the effect of photons promoting electrons to higher energy bands of the material, while at least one resonance in the infrared captures excitation of vibrational modes of the glass [44]. The refractive index of fused silica is shown in the top panel of Fig. 3, indicating both a UV resonance at the short wavelengths and the decrease in due to lattice absorption at long wavelengths.
@@ IV.3 Sources of Birefringence
When the core of the fiber is not circular, the boundary conditions that determine the propagation constants are different. Letting the ellipse be oriented with its semimajor and semiminor axes along the - and -axes (explicitly, ), we obtain two effective parameters . We also define the eccentricity 111If , all of the expressions in this section flip .. For small core eccentricities, we can use the Gaussian approximation to derive the resulting birefringence [100]. For and to second order in , it is [100, 73, 18]
@@
Here is the operating temperature of the fiber, , and are the dimensionless photoelastic constants of the core, and are the thermal expansion coefficients for the core and cladding, and is the dimensionless Poisson’s ratio for the core.
@@
where [dimensionless] is a commonly defined combination of material constants loosely called the “strain-optic coefficient.”
@@
When a fiber is twisted, a shear stress is introduced, which couples the transverse field of one mode to the longitudinal field of another. (Because we work in the weakly guiding limit, we have largely been ignoring the longitudinal field components, but in a typical Corning SMF-28e fiber at 1550 nm, the longitudinal field amplitude is a few percent of the transverse field amplitude.) The transverse and longitudinal fields are out of phase, so the coupling results in a circular birefringence. The birefringence is proportional to the twist rate [rad/m], and can be written [90]
@@ IV.5 Modeling Spun Fiber
where is the Pauli vector and is the identity matrix. The family of Jones matrices sampled in this way implements uniform rotations on the Poincaré sphere [32, 31].
@@ IV.6 Differential Group Delay
such that the total delay between these two modes is exactly from Eqn. 21. The eigenmodes in this case are and (where T is the transpose), which are sensible as these states are aligned with the fast and slow axes of the fiber. The eigenmodes are called principal states of polarization; they do not vary to first order with frequency, and the time delay between them is exactly the DGD of the time-domain picture.
@@
The DGD is a single scalar value, and therefore insufficient to describe the full PTF. Nevertheless, the DGD is still helpful as a heuristic metric of polarization rotation in the fiber. To see this, let us consider putting a temporally long pulse through each of three different fibers and consider the consequence of changing the fiber temperature. (1) If the fiber has no birefringence, then there is no PMD, and therefore zero DGD, and also the SOP remains the same at the output as at the input. The temperature change has no effect on the output SOP. (2) Conversely, in a long segment of fiber with constant birefringence, the delay between the two polarization components is allowed to build up over a long distance, so the total DGD is large. At the same time the SOP is also rotating along the fiber length, so the output SOP is very sensitive to changes in temperature. (3) Finally, for a long spun fiber, the spinning scrambles the birefringence axes; thus no two polarization modes are allowed to build up large relative delays, and the total DGD is small (in comparison to an unspun fiber of the same length). As the light traverses the fiber, the SOP does a random walk on the Poincaré sphere, rotating around one axis for a short length, then another axis for a short length, and so on. Such a random walk is hard to track, but the polarization does not wander as quickly as a function of length down the fiber. When the temperature is changed, each of these random walk steps is lengthened or shortened slightly, making it challenging to predict how the output SOP changes, but since the changes of length are in arbitrary directions, the total change to the SOP is smaller than for the unspun fiber of the same length.
@@ V Calculation Data and Methods
Our aim in this work is to model typical commercial telecommunications fiber links. Much of this fiber is predominantly silica-germania glass [91], i.e. pure silica (SiO2) for the cladding, and silica doped with a small amount of germania (GeO2) for the core, so we model these materials in our work. (Future extensions to BIFROST may include adding additional materials.) The material properties of bulk silica and bulk germania have been well-studied, giving us a wealth of data to use in BIFROST. To specify the effects of germania doping, we use the mole percent (%mol) (the fraction of molecules in the glass that are germania [91]) to specify the concentration, and we use a linear additive model for each of the relevant fiber properties. That is, for a given fiber property which has a value for pure silica and for pure germania, the value we use for silica with molar fraction of germania doping is
@@
To calculate indices of refraction, we use three-term Sellmeier equations for both silica and germania. For silica, we use the Sellmeier coefficients from Ref. [66]. The authors measured the refractive index of Corning 7980 fused silica in the ranges and , and fit the data to a set of three Sellmeier resonances with quartic temperature dependence. Thus the Sellmeier equation for silica takes into account the temperature dependence of the refractive index. For germania, we use the data from Refs. [40, 87], measured at an average of 24∘C over the wavelength range from 365 nm to 4.3 m. The authors do not know of any literature on the temperature-dependence of the coefficients of a three-term Sellmeier model of germania (though see Ref. [14] for the two-term case), but Ref. [91] gives an expression, fitted from measurements, for the thermo-optic coefficient of bulk germania glass in the range 200-300 K at 1550 nm. We therefore calculate the refractive index of germania at a temperature by calculating the expected value at 24∘C from Refs. [40, 87] and adding the expected change in the refractive index using the thermo-optic coefficient expression of Ref. [91].
@@ VI Validation of BIFROST
Having assembled a model of PMD in optical fibers, we need to verify that this model reproduces experimentally observed results. In some sense there is a wealth of data against which we can compare BIFROST results, since many works have made measurements of PMD in optical fibers (e.g. Refs. [61, 19, 16, 20] from the telecommunications engineering literature and Refs. [77, 64, 30] from the more recent quantum network literature). But in a more rigorous sense, the necessary experimental data for validation is scarce. BIFROST brings together many different “submodels,” including the model of the propagation constant of the fundamental modes Eqn. 2, the models of different birefringence mechanisms Eqns. 6, 8, 10, and 11, and the linear additive model for germania doping Eqn. 25 described in the last section. Ideally, each of these submodels should be compared to experiments which separate out the effects of the submodel under test, which is challenging experimentally. Such experiments, to the extent it is even possible to perform them, are rare in the literature. For this reason, validation of BIFROST is ongoing, and we are explicit about what we have examined thus far. We divide validation methodologies into five categories.
@@
In addition to verifying intrinsic fiber properties, it is important to verify the models of birefringence used in BIFROST. While all the equations we used are conventional models, each of them has limits; for instance, Eqn. 6 only works in the limit of small ellipticity, and Eqn. 10 may fail when the bend radius is small (i.e. when the bend is sharp). Thus it would be useful to check that each model works within reasonable regimes and test the degree of their failure in extreme regimes. To test such models requires very simple fiber systems whose properties we know well.
As an example of this type of validation, we consider fiber paddle systems such as ThorLabs FPC563. ThorLabs specifies [75] the total retardance of a single paddle given the paddle radius and the number of turns of fiber on the paddle: for instance, six turns of fiber with cladding diameter of m wound onto a 56 mm diameter paddle should yield a retardance of half a wave at nm. Fig. 5 confirms that BIFROST reproduces this result. We simulated a single paddle with the above parameters, and we specified input light with linear polarization at 1400 nm. The output polarization is shown on the Poincaré sphere as the paddle angle is varied, equivalent to turning a waveplate in free space 444We also specify in the simulation that there is 2 cm of straight fiber length at the input of the paddle, which twists as the paddle is rotated; we “measure” the polarization right out of the loop. The twist rotates the birefringence axes relative to the input, which makes it physically similar to rotating a waveplate in free space.. The resulting polarization rotation is almost exactly that expected of a rotating half-wave plate.
@@
As was discussed previously, there have been a number of previous attempts to model PMD in long fibers; some of these works ([e.g. 53, 20, 73, 18]) report on the specifics of their modeling choices and discuss specific birefringence mechanisms. In general, these works are not recent, and in the rare instances these works include comparison to experiment, the experiment is not well-documented. Nevertheless, comparisons between BIFROST and these previous results may show that BIFROST can reproduce previous computational results.
@@
In both cases, qualitative agreement is achieved: the birefringences due to core ellipticity and asymmetric thermal stress show similar shapes as a function of the normalized frequency , and the histograms from BIFROST fit well to Maxwellian curves just like the ones from Ref. [53]. However, there is quantitative disagreement of evident in both comparisons. We believe this discrepancy is mostly due to inconsistencies in Ref. [53] itself, not in BIFROST 555As a consistency check, we substituted Ref. [53]’s material values into Ref. [53]’s own expression for the asymmetric thermal stress (Eqn. 3), and the resulting birefringence values differ from their own Figure 2, which is reproduced in the upper left panel in our Fig. 6. The numbers we get this way vary from BIFROST’s numbers by under 10%, and the discrepancy is fully explained by slightly different values of material parameters. Given the internal numerical inconsistency in Ref. [53], we are inclined to support BIFROST’s calculations.. We conclude that the comparisons of Fig. 6 point to good agreement, adding to the validity of BIFROST.
@@
Lastly, we may try to use BIFROST to directly simulate fibers whose PMD values are measured in experiments. Candidates for this type of comparison have been identified [24, 61, 19, 77, 64]. In all of these cases, the fiber is not so well-reported that the simulation parameters to be used in BIFROST are apparent. However, BIFROST’s ability to analyze statistical ensembles allows us to explore the available parameter space and understand the parameter regimes under which BIFROST achieves qualitative and quantitative agreement with observations.